OpenStudy (sierra_tibbetts):

A system of equations is shown below: 5x + 2y = 3 (equation 1) 2x − 3y = 1 (equation 2) A student wants to prove that if equation 2 is kept unchanged and equation 1 is replaced with the sum of equation 1 and a multiple of equation 2, the solution to the new system of equations is the same as the solution to the original system of equations. If equation 2 is multiplied by 1, which of the following steps should the student use for the proof?

OpenStudy (sierra_tibbetts):

ANSWER CHOICES A) Show that the solution to the system of equations −4x + 9y = 5 and 2x − 3y = 1 is the same as the solution to the given system of equations B) Show that the solution to the system of equations 7x − y = 4 and 2x − 3y = 1 is the same as the solution to the given system of equations C) Show that the solution to the system of equations 2x + 5y = 3 and 3x − 2y = 1 is the same as the solution to the given system of equations D) Show that the solution to the system of equations 9x + 4y = 5 and 7x − y = 4 is the same as the solution to the given system of equations

OpenStudy (sierra_tibbetts):

@.Sam.

OpenStudy (sierra_tibbetts):

@3mar

OpenStudy (3mar):

Well, I am here.

OpenStudy (3mar):

Any ideas?

OpenStudy (3mar):

I am with you if you faced any difficulties! Let me know when you are back!